# IF-Critic: Towards a Fine-Grained LLM Critic for Instruction-Following Evaluation

### v1

@@ Appendix A Details of IF-Critic Training
We use the Llama-Factory (Zheng et al. 2024) framework and the Zero Redundancy Optimizer (ZeRO) (Rajbhandari et al. 2020) stage 3 from the Deepspeed library (Rasley et al. 2020) to train IF-Critic. All experiments are conducted on 8 H800 GPUs. We utilize the AdamW (Kingma and Ba 2014) optimizer with a weight decay of 0.1, and set the maximum sequence length to 8192 tokens during all the experiments.
@@
The overall API cost for critique training data collection is approximately 4000\$. During the supervised fine-tuning, the peak learning rate is set to 5e-6 with a 10% warmup ratio and a linear scheduler. The batch size is 16, and training is conducted for 2 epochs. During the constraint-level preference-optimization of IF-Critic, the peak learning rate is set to 1e-6 with a 10% warmup ratio and a cosine scheduler. The batch size is 128. Training utilizes a sigmoid loss function with a beta value of 0.1 and spans 1 epoch. Following previous works (Hou et al. 2024), an additional SFT loss is added to the chosen critique with a weight of 0.1.
Appendix B Details of Training Instructions Collection in Instruction-Following Optimization
@@
To ensure the diversity and complexity of constraints, we first construct a comprehensive and hierarchical constraint taxonomy, which consists of 6 primary categories (i.e., format constraints, content constraints, behavioral constraints, style constraints, role constraints, and environmental constraints) and 54 secondary categories. During constraint synthesis, we randomly sample 2 to 5 secondary constraint categories and require Deepseek-R1 to introduce several corresponding constraints into the seed instructions, thereby increasing their complexity. In contrast to previous work that only considers independent and atomic constraints (Jiang et al. 2023; Ren et al. 2025; Peng et al. 2025), we further ask Deepseek-R1 to compose multiple constraints according to the composition types proposed in ComlexBench (Wen et al. 2024), including: (1) And: Coordination between different constraints, (2) Chain: Sequential completion of constraints, (3) Selection: Conditional selection of constraints, and (4) The nested structure of the above composition types. In this way, our constraint synthesis can more comprehensively encompass the complex instruction types in real-world scenarios.
@@
In this scenario, it is notable that most of the hard constraints cannot be reliably verified using automatic verification codes alone, such as constraints targeting for specific segments of the response (e.g., Generate 10 titles, each no more than 8 words) and the composition of soft and hard constraints (e.g., Bold all adjectives that express feelings, such as **happy**), as it is challenging for verification code to accurately extract the corresponding segments of the response that should follow these constraints. We provide some examples of the seed instructions and the corresponding final constructed complex instructions in Table 5.
@@
Seed Instruction Write a conversation between a Chinese international student and a Japanese student who decide to try a new restaurant together. While they’re having their meal, one of them suggests that they go to see a movie afterward.
Final Instruction Please write a conversation between Xiaolin, a Chinese international student, and Yamada, a Japanese student. In the conversation, they decide to try a new restaurant together. While they’re having their meal, one of them suggests that they go to see a movie afterward. Ensure the following requirements are satisfied:
@@
3. Xiaolin should naturally mention Mid-Autumn Festival customs in the conversation, while Yamada should incorporate elements of the Obon Festival. Please use specific festival details (e.g., types of mooncakes or forms of Bon Odori).
@@
Appendix C Details of Instruction-Following Optimization
For DPO training, we use the Llama-Factory (Zheng et al. 2024) framework, and all other settings are the same as the preference optimization for IF-Critic. During the reward calculation, IF-Critic and QwQ-32B utilize the vllm (Kwon et al. 2023) inference framework, while Skywork-Reward-V2-Llama-3.1-8B-40M is deployed by sglang [^4] https://github.com/sgl-project/sglang (since this reward model is not compatible with the vllm framework currently). The reward calculation and DPO training are both conducted on 8 H800 GPUs.
For GRPO training, we use the VeRL (Sheng et al. 2025) framework. The learning rate is set to 2e-6, with the maximum sequence length for both prompts and responses set to 4096 tokens. The train-batch size and PPO mini-batch size are each set to 32. For rollout responses generation, we employ the vllm (Kwon et al. 2023) framework, utilizing 60% of the available GPU memory, with a temperature and top-p value of 1.0. To encourage stable training, we also incorporate KL divergence regularization with a coefficient of 0.001, using the low-variance KL implementation. We save checkpoints every 30 steps during training. Following TULU 3 (Lambert et al. 2024) and VerIF (Peng et al. 2025), we use IFEval (Zhou et al. 2023) as the validation set to select the best checkpoint. All experiments are conducted on 24 H800 GPUs, with 8 GPUs dedicated to model training and the remaining 16 GPUs allocated for API deployment of the instruction-following critic or reward model. The deployment framework is identical to the inference framework described in the above DPO training.
@@ Appendix C Details of Instruction-Following Optimization
Appendix D Reward Curves During GRPO Training
@@
Appendix E Experimental Results on General Benchmarks
We provide the performance of LLMs after instruction-following optimization in four general benchmarks: AlignBench (Liu et al. 2024), MMLU-Pro (Wang et al. 2024b), Omni-MATH (Gao et al. 2025), and HumanEval (Chen et al. 2021) in Table 6. GPT-4o-2024-11-20 is used as the evaluation model for AlignBench, and we use OmniJudge [^5] https://huggingface.co/KbsdJames/Omni-Judge for the evaluation of Omni-MATH. The results reveal that conducting instruction-following optimization with IF-Critic does not harm the general performance of LLMs.
Appendix F List of Prompt Templates
This section lists all the prompt templates applied throughout this work, including the prompt for scoring the quality of user instruction in Table 7, the prompt for constraint checklist generation in Table 8, the prompt for checklist-guided critique generation in Table 9, the prompt for cross-model verification in Table 10, and the prompt for rule-augmented verification in Table 11 and 12.
Appendix G Human Annotation Guideline for Constraint Verification
@@
1. Low Quality: The user instruction contains significant issues, such as unclear, incomplete, or ambiguous information, making it impossible to determine the intent behind the instruction.
@@
(2) You must extract and output all constraints in the order in which they appear in the instruction, without omitting any or inventing constraints not present in the instruction. To reiterate: Constraints refer to all instructional content, excluding auxiliary information such as background knowledge, text materials, and in-context examples.
@@
We will provide you with a user instruction, a constraint checklist for the user instruction, and an AI assistant’s response. Please think carefully and provide a detailed analysis of whether the AI assistant’s response follows each item on the constraint checklist. You need to analyze and assess every constraint in the checklist, and you do not need to analyze anything that is not listed in the checklist. You must output your explanation and judgment for each constraint, strictly following the format below:
@@
(2) Your judgment should be as strict as possible. You should output ”[[The AI assistant’s response follows this constraint]]” only if the response completely follows every part of the constraint as stated, without any omission or mistake.
@@
We will provide you with a user instruction, an AI assistant’s response, a specific constraint from the user instruction, and a critique of the AI’s response regarding whether the response follows this constraint from the user instruction. The critique comprises two parts: ”Explanation” and ”Judgment”. Your task is to carefully analyze this critique to judge whether it is reasonable and correct. After your analysis, output either ”[[The given critique is correct]]” or ”[[The given critique is not correct]]” as your final conclusion.
@@
(2) Be as strict as possible. Output ”[[The AI assistant’s response follows this requirement]]” only if the response completely follows every part of the requirement as stated, without any omission or mistake.
@@
(1) If the explanation section determines that the response follows the constraint, the judgment should be: ”[[The AI assistant’s response follows the constraint]]”.
(2) If the explanation section identifies any mistake or shortcoming in the response, indicating that the constraint is not fully followed, the judgment should be: ”[[The AI assistant’s response does not follow the constraint]]”.
@@
(1) There should not be situations such as ”partially follow the constraint”. As long as the explanation identifies any mistakes or shortcomings in the response that result in the constraint not being fully followed, the judgment should be: ”[[The AI assistant’s response does not follow the constraint]].”
(2) If the explanation states that the constraints and the response are unrelated and there is no need for evaluation, the judgment should be: ”[[The AI assistant’s response follows the constraint]].”
Please analyze the logical consistency between the explanation and judgment following the above principles. Output your analysis first, and then output either ”[[Explanation and Judgment are consistent]]” or ”[[Explanation and Judgment are not consistent]]” to represent your final verdict.
@@
We will provide you with a user instruction, an AI assistant’s response, and a specific constraint from the user instruction. Please first determine whether the given constraint is related to length (e.g., ”The article should be no less than 800 words” or ”Each title must be 8 characters long”). If so, please extract the segments from the AI assistant’s response that should follow this length requirement (extract the original text from the response without any modifications).
@@
”””json
   ”Length Constraint” : False
@@
”””json
   ”Length Constraint” : True,
   ”Extracted Segments” : [
       ”Length Requirement within the Constraint”: …(Provide the length requirement within the constraint. Be sure to extract the original text from the constraint without making any modification),
       ”Corresponding Segment in Response”: …(Provide the segment of the response that should follow this requirement without making any modification. If no corresponding segment exists in the response, output ”No corresponding segment exists.”)
       ”Length Requirement within the Constraint”: …,
       ”Corresponding Segment in Response”: …
@@
The given constraint is related to length. We will also provide a JSON data, which contains multiple items. Each item consists of three elements: (1) a length requirement within the constraint, (2) a segment of the responses that should follow the length requirement, and (3) the actual length of the segment. It is important that you consider the ”actual length provided in the JSON data” as the golden-truth length, which is absolutely accurate. Do not make your own calculation about “the length of the segment”. You must rely solely on the ”actual length” as the basis for length validation. Output your explanation and judgment for the given constraint based on the actual length provided in the JSON data, strictly following the format below:
@@
(2) Your judgment should be as strict as possible. You should output ”[[The AI assistant’s response follows this constraint]]” only if the response completely follows every part of the constraint as stated, without any omission or mistake.
(3) Independently analyzing the given constraint, without taking into account whether other constraints are followed.
(4) The given constraint may contain both length requirements and other requirements. For length requirements, you must rely solely on the ”actual length provided in the JSON data” as the basis for length validation. For other requirements, please refer to the above principles and carefully evaluate them by yourself.
(5) When generating your explanation, it is forbidden to mention phrases such as ”according to the given JSON data” or ”based on the provided actual length”. Your explanation should convey that the length is calculated independently by you, without referencing the given JSON data.
@@
Below, an instruction, a corresponding model response, and a constraint checklist for the instruction will be provided. The checklist contains some of the constraints within the instruction. Your task is to verify whether the model response follows each constraint in the checklist, choose ”Followed” or ”Not Followed”, and provide a brief justification.
@@
2. Your judgments must be strict. Only responses that fully satisfy a constraint can be judged as ”Followed”. If there is any omission or error regarding a constraint, it must be judged as ”Not Followed”.

### v2

@@ Appendix D Implementation Details of IF-Critic
We use the Llama-Factory Zheng et al. (2024) framework and the Zero Redundancy Optimizer (ZeRO) Rajbhandari et al. (2020) stage 3 from the Deepspeed library Rasley et al. (2020) to train IF-Critic. All experiments are conducted on 8 H800 GPUs. We utilize the AdamW Kingma and Ba (2014) optimizer with a weight decay of 0.1, and set the maximum sequence length to 8192 tokens during all the experiments.
@@
During the supervised fine-tuning, the peak learning rate is set to 5e-6 with a 10% warmup ratio and a linear scheduler. The batch size is 16, and training is conducted for 2 epochs. During the constraint-level preference-optimization of IF-Critic, the peak learning rate is set to 1e-6 with a 10% warmup ratio and a cosine scheduler. The batch size is 128. Training utilizes a sigmoid loss function with a beta value of 0.1 and spans 1 epoch. Following previous works Hou et al. (2024), an additional SFT loss is added to the chosen critique with a weight of 0.1.
The overall API cost for critique training data collection is approximately 4000\$. During final explanation selection, we use difflab [^2] https://docs.python.org/3/library/difflib.html to implement the text similarity measure $u(\cdot)$.
Appendix E Experimental Results on Other Base Models
Base Model EvalCritic CFBench TRACE Multi-IF
Qwen2.5-14B-Insturct 0.861 0.863 0.840 0.895
Qwen-2.5-7B-Instruct 0.822 0.798 0.806 0.813
Qwen-3-8B 0.855 0.823 0.820 0.848
GLM-4-9B-0414 0.797 0.810 0.801 0.810
Qwen-3-14B 0.868 0.863 0.828 0.860
Qwen-2.5-32B-Instruct 0.879 0.854 0.833 0.867
Table 5: The average of Positive F1 and Negative F1 scores of IF-Critic under different base models. Qwen-2.5-14B-Instruct is the base model we used for the main experiments.
In our primary experiment, we experiment with various base models for IF-Critic, and the results are presented in Table 5. We observe that LLMs with approximately 8B parameters are still insufficient to fully acquire critique generation capabilities, exhibiting a notable performance gap to larger LLMs. In contrast, LLMs with approximately 14B parameters are generally adequate for learning critique generation, and further increasing model scale yields negligible performance improvement. Considering the trade-off between performance and efficiency, we ultimately choose Qwen-2.5-14B-Instruct as the base model for our main experiments.
Appendix F Details of Instruction-Following Optimization
Training Instructions Collection.
@@ Appendix F Details of Instruction-Following Optimization
To ensure the diversity and complexity of constraints, we first construct a comprehensive and hierarchical constraint taxonomy, which consists of 6 primary categories (i.e., format constraints, content constraints, behavioral constraints, style constraints, role constraints, and environmental constraints) and 54 secondary categories. During constraint synthesis, we randomly sample 2 to 5 secondary constraint categories and require Deepseek-R1 to introduce several corresponding constraints into the seed instructions, thereby increasing their complexity. In contrast to previous work that only considers independent and atomic constraints Jiang et al. (2023); Ren et al. (2025); Peng et al. (2025), we further ask Deepseek-R1 to compose multiple constraints according to the composition types proposed in ComlexBench Wen et al. (2024), including: (1) And: Coordination between different constraints, (2) Chain: Sequential completion of constraints, (3) Selection: Conditional selection of constraints, and (4) The nested structure of the above composition types. In this way, our constraint synthesis can more comprehensively encompass the complex instruction types in real-world scenarios.
@@
In this scenario, it is notable that most of the hard constraints cannot be reliably verified using automatic verification codes alone, such as constraints targeting for specific segments of the response (e.g., Generate 10 titles, each no more than 8 words) and the composition of soft and hard constraints (e.g., Bold all adjectives that express feelings, such as happy), as it is challenging for verification code to accurately extract the corresponding segments of the response that should follow these constraints. We provide some examples of the seed instructions and the corresponding final constructed complex instructions in Table 6.
@@
Seed Instruction Write a conversation between a Chinese international student and a Japanese student who decide to try a new restaurant together. While they’re having their meal, one of them suggests that they go to see a movie afterwards.
Final Instruction Please write a conversation between Xiaolin, a Chinese international student, and Yamada, a Japanese student. In the conversation, they decide to try a new restaurant together. While they’re having their meal, one of them suggests that they go to see a movie afterwards. Ensure the following requirements are satisfied:
@@
3. Xiaolin should naturally mention Mid-Autumn Festival customs in the conversation while Yamada should incorporate elements of the Obon Festival. Please use specific festival details (e.g., types of mooncakes, or forms of Bon Odori).
@@
Training Configuration.
For DPO training, we use the Llama-Factory Zheng et al. (2024) framework, and all other settings are the same as the preference optimization for IF-Critic. During the reward calculation, IF-Critic and QwQ-32B utilize the vllm Kwon et al. (2023) inference framework, while Skywork-Reward-V2-Llama-3.1-8B-40M is deployed by sglang [^3] https://github.com/sgl-project/sglang (since this reward model is not compatible with the vllm framework currently). The reward calculation and DPO training are both conducted on 8 H800 GPUs.
For GRPO training, we use the VeRL Sheng et al. (2025) framework. The learning rate is set to 2e-6, with the maximum sequence length for both prompts and responses set to 4096 tokens. The train-batch size and PPO mini-batch size are each set to 32. For rollout responses generation, we employ the vllm Kwon et al. (2023) framework, utilizing 60% of the available GPU memory, with a temperature and top-p value of 1.0. To encourage stable training, we also incorporate KL divergence regularization with a coefficient of 0.001, using the low-variance KL implementation. We save checkpoints every 30 steps during training. Following TULU 3 Lambert et al. (2024) and VerIF Peng et al. (2025), we use IFEval Zhou et al. (2023) as the validation set to select the best checkpoint. All experiments are conducted on 24 H800 GPUs, with 8 GPUs dedicated to model training and the remaining 16 GPUs allocated for API deployment of the instruction-following critic or reward model. The deployment framework is identical to the inference framework described in the above DPO training.
@@
Appendix G Reward Curves During GRPO Training
@@
Appendix H Experimental Results on General Benchmarks
We provide the performance of LLMs after instruction-following optimization in four general benchmarks: AlignBench Liu et al. (2024), MMLU-Pro Wang et al. (2024), Omni-MATH Gao et al. (2025), and HumanEval Chen et al. (2021) in Table 7. GPT-4o-2024-11-20 is used as the evaluation model for AlignBench, and we use OmniJudge [^4] https://huggingface.co/KbsdJames/Omni-Judge for the evaluation of Omni-MATH. The results reveal that conducting instruction-following optimization with IF-Critic does not harm the general performance of LLMs.
Appendix I Human Annotation Guideline for Constraint Verification
@@
1. Low Quality: The user instruction contains significant issues such as unclear, incomplete, or ambiguous information, making it impossible to determine the intent behind the instruction.
@@
(2) You must extract and output all constraints in the order in which they appear in the instruction, without omitting any constraints or inventing constraints not present in the instruction. To reiterate: Constraints refer to all instructional content, excluding auxiliary information such as background knowledge, text materials, and in-context examples.
@@
We will provide you with a user instruction, a constraint checklist for the user instruction, and an AI assistant’s response. Please think carefully and provide a detailed analysis of whether the AI assistant’s response follows each item on the constraint checklist. You need to analyze and assess every constraint in the checklist, and do not need to analyze anything that is not listed in the checklist. You must output your explanation and judgment for each constraint, strictly following the format below:
@@
(2) Your judgment should be as strict as possible. You should output "[[The AI assistant’s response follows this constraint]]" only if the response completely follows every part of the constraint as stated, without any omission or mistake.
@@
We will provide you with a user instruction, an AI assistant’s response, a specific constraint from the user instruction, and a critique of the AI’s response regarding whether the response follows this constraint from the user instruction. The critique comprises two parts: "Explanation" and "Judgment". Your task is to carefully analyze this critique to judge whether it is reasonable and correct. After your analysis, output either "[[The given critique is correct]]" or "[[The given critique is not correct]]" as your final conclusion.
@@
(2) Be as strict as possible. Output "[[The AI assistant’s response follows this requirement]]" only if the response completely follows every part of the requirement as stated, without any omission or mistake.
@@
(1) If the explanation section determines that the response follows the constraint, the judgment should be: "[[The AI assistant’s response follows the constraint]]".
(2) If the explanation section identifies any mistake or shortcoming in the response, indicating that the constraint is not fully followed, the judgment should be: "[[The AI assistant’s response does not follow the constraint]]".
@@
(1) There should not be situations such as "partially follow the constraint". As long as the explanation identifies any mistakes or shortcomings in the response that result in the constraint not being fully followed, the judgment should be: "[[The AI assistant’s response does not follow the constraint]]."
(2) If the explanation states that the constraints and the response are unrelated and there is no need for evaluation, the judgment should be: "[[The AI assistant’s response follows the constraint]]."
Please analyze the logical consistency between the explanation and judgment following the above principles. Output your analysis first, and then output either "[[Explanation and Judgment are consistent]]" or "[[Explanation and Judgment are not consistent]]" to represent your final verdict.
@@
We will provide you with a user instruction, an AI assistant’s response, and a specific constraint from the user instruction. Please first determine whether the given constraint is related to length (e.g., "The article should be no less than 800 words" or "Each title must be 8 characters long"). If so, please extract the segments from the AI assistant’s response that should follow this length requirement (extract the original text from the response without any modifications).
@@
"""json
   "Length Constraint" : False
@@
"""json
   "Length Constraint" : True,
   "Extracted Segments" : [
       "Length Requirement within the Constraint": …(Provide the length requirement within the constraint. Be sure to extract the original text from the constraint without making any modification),
       "Corresponding Segment in Response": …(Provide the segment of the response that should follow this requirement without making any modification. If no corresponding segment exists in the response, output "No corresponding segment exists.")
       "Length Requirement within the Constraint": …,
       "Corresponding Segment in Response": …
@@
The given constraint is related to length. We will also provide a JSON data, which contains multiple items. Each item consists of three elements: (1) a length requirement within the constraint, (2) a segment of the responses that should follow the length requirement, and (3) the actual length of the segment. It is important that you consider the "actual length provided in the JSON data" as the golden-truth length, which is absolutely accurate. Do not make your own calculation about “the length of the segment”. You must rely solely on the "actual length" as the basis for length validation. Output your explanation and judgment for the given constraint based on the actual length provided in the JSON data, strictly following the format below:
@@
(2) Your judgment should be as strict as possible. You should output "[[The AI assistant’s response follows this constraint]]" only if the response completely follows every part of the constraint as stated, without any omission or mistake.
(3) Independently analyzing the given constraint, and does not take into account whether other constraints are followed.
(4) The given constraint may contain both length requirements and other requirements. For length requirements, you must rely solely on the "actual length provided in the JSON data" as the basis for length validation. For other requirements, please refer to the above principles and carefully evaluate them by yourself.
(5) When generating your explanation, it is forbidden to mention phrases such as "according to the given JSON data" or "based on the provided actual length". Your explanation should convey that the length is calculated independently by you, without referencing the given JSON data.
@@
Below, an instruction, a corresponding model response, and a constraint checklist for the instruction will be provided. The checklist contains some of the constraints within the instruction. Your task is to verify whether the model response follows each constraint in the checklist, choose "Followed" or "Not Followed", and provide a brief justification.
@@
2. Your judgments must be strict. Only responses that fully satisfy a constraint can be judged as "Followed". If there is any omission or error regarding a constraint, it must be judged as "Not Followed".